HETAV PATEL

Grade 11,

if 1+sin a =9 cos a then 5(1-cos a) if 1+sin a =9 cos a & 0

if 1+sin a =9 cos a then 5(1-cos a)
if 1+sin a =9 cos a & 0

Grade:11

2 Answers

Vijay Mukati
7 years ago
Dear Student,

Please write your question clearly. Its not clearly formatted.

Thanks, Vijay
Vikas TU
7 years ago
Dear Yathisth,
as this question is uncleared i have answered this question.
check again,
1+sin a=9 cos a 
Divide by cosa both sides, we get
seca + tana = 9
seca = 9 – tana
squaring both sides,
(seca)^2 = 81 + (tana)^2  – 18tana
1 + (tana)^2  = 8` + tana^2 – 18tana
tan a = 7/18
cosa = 18/root(49 + 18^2)
sina = 7/root(49 + 18^2)
now cross check with all the options.